the amount of 8000 rupees is invested in bank , some at 6.5% and the rest at 4.5% How much should be invested in each in order to get an annual interest of rupees 420?
Answer by mananth(16946)

Fund A 6.50% per annum ---x
Fund B 4.50% per annum ---y

x + y= 8000 ------------------------1
6.50% x + 4.50% y= = 420.00
Multiply by 100
6.5 x + 4.5 y= = 42,000.00 --------2

Multiply (1) by -6.5
we get

-6.5 x -6.5 y= = -52000.00

Add this to (2)

0 x -2 y= = -10,000.00

divide by -2

y = 5,000.00 investment at 4.50%
Balance 3,000.00 investment at 6.50%
CHECK
$3,000.00 --------- 6.50% -------195.00
$5,000.00 ------- 4.50% ------- 225.00
Total -------------- 420.00